OpenStudy (anonymous):

using the slope-form how do i write an equation for a line parallel to 2x-3y=7 and passes though (5,7)

OpenStudy (anonymous):

slope of the original line is = 2/3 now to write the new equation, u use the point (5,7) and the equation y - y1 = m(x - x1), that is: y-7 = 2/3(x-5) y-7=2/3x-10/3 simplifying I got 2 x - 3 y - 31 = 0

OpenStudy (anonymous):

any question, please ask

OpenStudy (anonymous):

no questions i understand

OpenStudy (anonymous):

ok